Thursday, 20 August

Thursday, 20 August

Plenary Lectures

9:30 – 10:30  Auditorium Maximum H 105

Dusa McDuff, SUNY, Stony Brook, USA Fibrations in Symplectic Topology (I: p.16)

Chair: Yakov Eliashberg, Stanford University, Stanford, USA

11:00 – 12:00  Auditorium Maximum H 105

I.G. Macdonald, Queen Mary and Westfield College, University of London, London, England Constant Term Identities, Orthogonal Polynomials, and Affine Hecke Algebras (I: p.12)

Chair: Richard Borcherds, University of Cambridge, Cambridge, UK

Invited Lectures

\underline{14:00 – 14:45}

04 Mark L. Green, UCLA, Los Angeles, USA Higher Abel-Jacobi Maps (I: p.62) Chair: Frank Schreyer

MA 005

07 Alexander Eskin, University of Chicago, Chicago, USA Counting problems and semisimple groups (I: p.91) Chair: Peter Sarnak

MA 004

08 Sergey Ivanovich Pinchuk, Indiana University, Bloomington, USA and Chelyabinsk State Univ. of Technology, Chelyabinsk, Russia Reflection Principle in Higher Dimensions (I: p.104) Chair: Jürgen Leiterer

MA 001

10 Matei Machedon, University of Maryland, College Park, USA Fourier Analysis of Null Forms and Non-linear Wave Equations (I: p.124) Chair: Michael Struwe

H 1012

12 Ruth J. Williams, University of California at San Diego, La Jolla, USA Reflecting Diffusions and Queueing Networks (I: p.152) Chair: Ludwig Arnold

H 1028

17 Alexander Schrijver, CWI and University of Amsterdam, Amsterdam, Netherlands Routing and timetabling by topological search (I: p.194) Chair: Günter M. Ziegler

H 2032

\underline{15:00 – 15:45}

04 Maurizio Cornalba, Università di Pavia, Pavia, Italy Cohomology of moduli spaces of stable curves (I: p.61) Chair: Alberto Conte

MA 005

07 Joseph Bernstein, Tel Aviv University, Tel Aviv, Israel Analytic structures on representation spaces of reductive groups (I: p.88) Chair: Wolfgang Soergel

MA 004

08 Nigel David Higson, Pennsylvania State University, University Park, USA The Baum-Connes Conjecture (I: p.98) Chair: Joachim Cuntz

MA 001

10 Frank Eric Merle, Université de Cergy-Pontoise, Cergy-Pontoise, France Blow-up Phenomenon For Critical Nonlinear Schrödinger and Zakharov Equations

(I: p.124)

Chair: Elmar Schrohe H 1012 12 Maury D. Bramson, University of Minnesota, Minneapolis, USA State Space Collapse for Queueing Networks (I: p.141) Chair: Jennifer T. Chayes

H 1028

17 Michel X. Goemans, CORE, Louvain-La-Neuve, Belgium and MIT, Cambridge, USA Semidefinite programming and combinatorial optimization (I: p.193) Chair: András Frank

H 2032

\underline{16:15 – 17:00}

!!!

Lecture by a Fields Medalist

– to be announced – Chair: N.N.

MA 005

05 Joachim Lohkamp, Universität Augsburg, Augsburg, Germany Curvature Contents of Geometric Spaces (I: p.75) Chair: Jean-Pierre Bourguignon

MA 001

06 Tomotada Ohtsuki, Tokyo Institute of Technology, Tokyo, Japan A filtration of the set of integral homology 3-spheres (I: p.82) Chair: W.B. Raymond Lickorish

MA 004

09 Krystyna M. Kuperberg, Auburn University, Auburn, USA Counterexamples to the Seifert Conjecture (I: p.117) Chair: Urs Kirchgraber

H 1012

11 Jennifer T. Chayes, Microsoft Research, Redmond, USA Birth of the Infinite Cluster: Finite-Size Scaling in Percolation (I: p.130) Chair: Jean-Dominique Deuschel

H 1028

14 Madhu Sudan, MIT, Cambridge, USA Probabilistic checking of proofs (I: p.173) Chair: Hans J. Prömel

H 2032

\underline{17:15 – 18:00}

05 Yakov Eliashberg, Stanford University, Stanford, USA Invariants in Contact Topology (I: p.70) Chair: Jean-Pierre Bourguignon

MA 001

06 Clifford Henry Taubes, Harvard University, Cambridge, USA The Geometry of the Seiberg-Witten Invariants (I: p.86) Chair: Dieter Kotschick

MA 004

09 Håkan Eliasson, Royal Institute of Technology, Stockholm, Sweden Reducibility and point spectrum for linear quasi-periodic skew-products (I: p.112) Chair: Reiner Lauterbach

H 1012

11 Detlev Buchholz, Universität Göttingen, Göttingen, Germany Scaling algebras in local relativistic quantum physics (I: p.129) Chair: Gian Michele Graf

H 1028

14 Johan Håstad, Royal Institute of Technology, Stockholm, Sweden On approximating NP-hard optimization problems (I: p.169) Chair: Hans J. Prömel

H 2032

16 Graeme Walter Milton, The University of Utah, Salt Lake City, USA Exact relations for composites: towards a complete solution (I: p.191) Chair: Krzysztof Wilmanski MA 005

Short Communications and Poster Sessions

Short Communications: Thursday, 20 August, 14:00–14:45

14:00–14:15  14:15–14:30  14:30–14:45

Sect.: 01

Lattices and Universal Algebra, Chair: Marcus A. Kracht

MA 042 Isabella C. Burger Boolean algebras and informationdependent entailment relations

(p.6)

Dietmar Dorninger

On ring-like structures occurring in axiomatic quantum mechanics (p.7)

Philippe Niederkorn

Natural dualities for finitely generated varieties of MV-algebras (p.8)

Sect.: 03

Cyclotomy, Chair: Stanley J. Gurak

MA 311 Grzegorz, M. Banaszak Wild kernel for higher K-groups and arithmetic of square rings (p.34)

Mónica del Pilar Canales

Higher levels of Z/p^{l}Z and cyclotomy (p.34)

Stanley J. Gurak

On the factors Φ^{(jδ/m)}(x) of the period polynomial for finite fields (p.37)

Sect.: 05

Special Submanifolds I, Chair: Takashi Okayasu

MA 041 Vladimir G. Tkachev Upper estimates of volume growth for minimal quasigraphs (p.82)

Liu Ximin

Positively curved totally complex submanifolds in the Cayley projective space (p.75)

Takashi Okayasu

A Bernstein result for minimal graphs of higher codimension (p.77)

Sect.: 06

Gauge Theory and Low Dimensional Topology, Chair: Ronald J. Stern

H 112 Yong S. Cho Finite Group Actions and Seiberg-Witten Invariants (p.87)

Nikolai Saveliev

Casson invariant and surface singularities (p.97)

Louis H. Kauffmann

Functional Integrals and the Kontsevich Integral (p.90)

Sect.: 09

Chaotic Dynamical Systems, Chair: Mohamed Houssni

H 111 Abdelhak Fahsi Saddle loop and chaotic bifurcations in nonlinear oscillators (p.175)

Jaroslav Smital

Distributional chaos in dynamical systems generated by iterates of continuous maps

(p.192)

Mohamed Houssni

Suppression of chaos in a quasi-periodically excited system (p.177)

Sect.: 09

Periodic Solutions of Differential Equations, Chair: Mohammad Jahanshahi

H 110 Mihail Emilian Popescu Existence and Oribital Stability of Periodic Solutions for Small-Parameter Non-Linear Differential Systems (p.186)

Sergiy V. Yanchuk

Conditions for the existence of cycles in some two-dimensional system of differential equations (p.195)

Mohammad Jahanshahi

Periodic Solution of a Non-Linear Dynamical System of Second Order with Dimension n (p.178)

Sect.: 11

Integrable Systems I, Chair: Renat Z. Zhdanov

H 107 Igor Korepanov Integrability in higher dimensions (p.238)

Iordan P. Michev

On the Fay identity for Kdv tau functions and the identity for the Wronskian of squared solutions of Sturm-Liouville equation (p.240)

Renat Z. Zhdanov

On classification of separable (2+1)- dimensional Schrödinger equations

(p.246)

Sect.: 14

Mathematical Aspects of Computer Science I, Chair: Stefan Felsner

MA 043 Laxminarayan Das Convexity to Invexity (p.290)

Narayan C. Debnath

RGPG: A Graph for Software Modeling (p.290)

Victor Kasyanov

Hierarchical graphs and visual processing (p.292)

Sect.: 15

Approximation I, Chair: Jozsef Grof

H 2033 Alexis Bacopoulos On the link among vectorial minimal projections and best approximations (p.296)

Igor Fedotov

Another Approach to Quadrature Methods (p.300)

Jozsef Grof

Einige Verallgemeinerungen bzw. Modifizierungen der Szász-Mirakjan-Operatoren (p.301)

Sect.: 15

Partial Differential Equations I, Chair: Vladimir V. Shaidurov

H 2013 Christoph Pflaum Robust Multilevel Algorithms for Elliptic Differential Equations (p.309)

Vladimir V. Shaidurov

Cascade finite element methods for weakly nonlinear elliptic equations (p.313)

Urve Kangro

Spectral Convergence of the Interior Source Method with Point Matching

(p.303)

Sect.: 16

Non-linear Wave Propagation, Chair: Lokenath Debnath

H 106 Mikhail Vasiliev Automodelling solutions to the equations of the viscous heat conductive gas flows

(p.333)

Mark C.W. Jones

Coupled evolution equations and the stability of stratified flows (p.325)

Lokenath Debnath

Resonant Phenomena in Linear and Nonlinear Water Waves (p.321)

Short Communications: Thursday, 20 August, 15:00–15:45

15:00–15:15  15:15–15:30  15:30–15:45

Sect.: 03

Function Fields and Curves I, Chair: Franck Leprévost

MA 311 Florian Heß Effective construction of Riemann-Roch spaces in global function fields (p.38)

Ryutaro Horiuchi

On Weierstrass Points of the Fermat Curve (p.38)

Franck Leprévost

On some elliptic surfaces and elliptic curves related to discriminants of cubic or quartic polynomials (p.41)

Sect.: 05

Special Submanifolds II, Chair: Ivko M. Dimitric

MA 041 Valentin Gabriel Cristea Totally umbilical graded submanifolds

(p.66)

S.M.B. Kashani

Cohomogeneity One Revolution Hypersurfaces of the Euclidean Space (p.72)

Ivko M. Dimitric

Low-type submanifolds of non-flat complex space forms (p.67)

Sect.: 06

Knots and 3-Manifolds, Chair: Jonathan A. Hillman

H 112 Joel Hass The number of Reidemeister Moves Needed for Unknotting (p.89)

Cherry Kearton

Alexander ideals of classical knots (p.90)

Michelangelo Vaccaro

Unicity of the topological type of the simply connected 3-manifolds (p.99)

Sect.: 09

Boundary Value Problems Related to FDE, Chair: Zhanyuan Hou

H 111 Manuel J. Alves

�  �������  �  ������������
�����������˘�  ������˘�  ������

��� ������������ �����������- ��-����������������� ���������

(p.169)

Sergei A. Brykalov

Solution sets of problems with nonlinear functional boundary conditions (p.173)

Zhanyuan Hou

L^{p}-Perturbation of Linear Functional Differential Equations (p.176)

Sect.: 09

Periodic Solutions of FDE, Chair: Mustapha Yebdri

H 110 Nataliya Bantsur Existence and stability of the periodic and almost periodic solutions of differential equations with maxima (p.171)

Joao S.P. Munembe

Computationally verifiable conditions of the existence of stabilizable solutions for linear differential systems with periodic parameters (p.184)

Mustapha Yebdri

The linear stability of slowly oscillating periodic solutions of a delay differential equation (p.195)

Sect.: 11

Statistical Mechanics, Chair: Ali E Özlük

H 107 Daniel Marinov Danchev FSS Properties of an Exactly Solvable Quantum Model (p.231)

Valentina Goloubeva

On algebraic similarity of Calogero and Sutherland models (p.234)

Ali E Özlük

A Farey Fraction Spin Chain (p.242)

Sect.: 14

Mathematical Aspects of Computer Science II, Chair: Clemens Gröpl

MA 043 Guy Merlin Mbakop An Efficient Algorithm for Solving Real Polynomial Equations: The Complete Intersection Case (p.294)

Bernard Mourrain

Multivariate polynomial equations and duality (p.294)

Emil J Volcheck

Computing with algebraic curves: a survey of recent results (p.295)

Sect.: 15

Approximation II, Chair: Alexei A. Mailybaev

H 2033 Alois Heinz Efficient function approximation by treestructured neural networks (p.302)

Peter Kirrinnis

Fast algorithms for partial fractions (p.305)

Alexei A. Mailybaev

Transformation of Families of Matrices to Normal Forms (p.306)

Sect.: 15

Partial Differential Equations II, Chair: Jürgen Fuhrmann

H 2013 Bodo Erdmann Mesh Controlling Strategy for Solving Nonlinear Evolution Problems (p.299)

Christian Wieners

Adaptive multigrid methods for Prandtl-Reuß-plasticity (p.318)

Jürgen Fuhrmann

Numerically Stable Finite Volume Discretizations for Nonlinear Parabolic Problems (p.300)

Sect.: 16

Geophysical and Environmental Models, Chair: Jinqiao Duan

H 106 Verdiana Grace Masanja A Mathematical Model and Simulation of Transport of Dust Particles in Dar es Salaam City (p.329)

Vitaly G. Il’ichev

Delta-functions and analysis of ecological interactions in a variable environment

(p.324)

Jinqiao Duan

Dynamical Behavior of Quasigeostrophic Models (p.322)

Short Communications: Thursday, 20 August, 16:15–17:00

16:15–16:30  16:30–16:45  16:45–17:00

Sect.: 02

Noetherian Rings, Chair: Kurt Rosenbaum

MA 311 Vijay Kodiyalam Asymptotic behaviour of Castelnuovo-Mumford regularity (p.20)

Syed Tariq Rizvi

On noetherian and artinian rings (p.26)

Sylvia M. Wiegand

Building Noetherian domains using completions (p.31)

Sect.: 04

Toric Varieties, Chair: Klaus Altmann

MA 042 Anette A’Campo-Neuen Quasi–Projective Reduction of Toric Varieties (p.49)

Jürgen Hausen

Toric Quotients (p.53)

Evgeni Nikolaevich Materov

One estimate for dimension of H^{n}(C^{n}\V, C) (p.56)

Sect.: 07

Symmetric Spaces, Chair: Walter J. Michaelis

H 112 Sergei S. Platonov Invariant subspaces of some function spaces on symmetric spaces (p.109)

Shigeru Aoki

Invariant Eigendistributions on U(1, 2)/(U(1, 1) × U(1, 1)) (p.101)

Adam Koranyi

Equivariant differential operators on homogeneous vector bundles over G/MAN

(p.107)

Sect.: 08

Sobolev- and Function Spaces, Chair: Michal Wojciechowky

H 110 Ritva Hurri-Syrjanen Generalizations of Poincaré inequality

(p.131)

Evgueni M. Semenov

Disjoint Systems and Complemented Subspaces in Rearrangement Invariant Spaces

(p.155)

Michal Wojciechowky

The multidimensional Marcinkiewicz multiplier theorem for H^{1} spaces on product domains (p.164)

Sect.: 08

Integral Transforms I, Chair: Vakhtang M. Kokilashvili

MA 041 Jagdish N. Pandey The Hilbert Transform of Almost Periodic Functions and Distributions (p.146)

Alexander N. Meskhi

Solution of some weight problems for the Riemann-Liouville and Weyl operators defined on SHT (p.141)

Vakhtang M. Kokilashvili

Solutions of Two-Weight Problems on Boundedness and Compactness for Integral Transforms (p.135)

Sect.: 08

Variational Problems, Chair: Julian P. Revalski

MA 309 Hong-Kun, H.K. Xu Asymptotic behavior of a gradient flow

(p.164)

Mikhail A. Sytchev

Characterization of homogeneous scalar variational problems solvable for all boundary data (p.158)

Julian P. Revalski

Porosity and variational principles (p.152)

Sect.: 08

Topological Vector Spaces and Algebras, Chair: Wieslaw Zelazko

MA 043 Aboubakr Bayoumi Quasi-Differentiability and Applications in Locally Bounded Spaces (p.116)

Mati Abel

Structure of locally pseudoconvex algebras (p.112)

Wieslaw Zelazko

A supersimple algebra (p.166)

Sect.: 10

Systems I, Chair: Yakov Krasnov

H 111 Gennadii V. Demidenko Systems not of Cauchy–Kovalevskaya type

(p.201)

Abduhamid Dzhuraev

On partial differential equations in the unit ball of ℂ^{n}, which are singular at the origin and on the boundary (p.202)

Yakov Krasnov

Symmetries of Cauchy-Riemann-Füter Equation (p.211)

Sect.: 12

Interacting Particle Systems, Chair: Hans-Otto Georgii

H 107 Jochen Geiger The genealogy of a cluster in the multitype voter model (p.251)

Alejandro F. Ramirez

Diffusion and saturation processes in random media (p.267)

Kenji Handa

A lower bound for time correlation of lattice gas (p.252)

Sect.: 12

Limit Theorems, Chair: Ludwig Paditz

H 1029 Alexander V. Bulinski Limit laws for arrays (p.249)

Oleg I. Klesov

A general form of the strong law of large numbers (p.256)

Michael P. Voit

Ultraspherical expansions and central limit theorems for unit spheres (p.271)

Sect.: 13

Combinatorics of Words, Chair: Emo Welzl

H 106 Julien Cassaigne On the conjectured subword complexity of the Kolakoski sequence (p.275)

Martin Klazar

Combinatorics of a containment relation that generalizes Davenport–Schinzel sequences (p.279)

Alexander Sergeevich Makarenko

The complexity measure of individual geometrical and algebraical objects (p.281)

Sect.: 15

Wavelets II (organized by Andreas Rathsfeld), Chair: Jacques Liandrat

H 2013 Rudolph A. Lorentz On the nonexistence of compactly supported Sobolev prewavelets for box spline multiresolution analyses (p.306)

Gunther Schmidt

Representation of multidimensional potentials by approximate wavelets (p.312)

Wayne M. Lawton

Energy functions on graphs, wavelets, and multilevel algorithms (p.305)

Sect.: 17

Nonlinear Optimization I, Chair: Jürgen Guddat

H 2033 Leonid Faybusovich Interior-point algorithms and Euclidean Jordan algebras (p.341)

Bernd Kummer

Metric Regularity in Optimization: Trivial and Nontrivial Situations (p.344)

Ider Tsevendorj

Sufficient Global Optimality Conditions for Nonconvex Nonsmooth Problems (p.353)

Short Communications: Thursday, 20 August, 17:15–18:00

17:15–17:30  17:30–17:45  17:45–18:00

Sect.: 02

Rings and Modules I, Chair: Sylvia M. Wiegand

MA 311 Toma Albu Global Krull dimension and global dual Krull dimension of rings (p.12)

Jilyana Cazaran

Sufficient conditions for finite commutative rings to be principal ideal rings and an application to error-correcting codes

(p.14)

Reinhard Börger

Infinitary Linear Combinations in Cotorsion Modules (p.14)

Sect.: 04

Singularities I, Chair: Gert-Martin W. Greuel

MA 042 Sabir M. Gusein-Zade On the index of a gradient vector field at an isolated complete intersection singularity

(p.53)

Ludwig Balke

Open book fibrations of lens spaces and deformations of cyclic quotient singularities

(p.50)

Claus Hertling

Variational Torelli result for hypersurface singularities (p.53)

Sect.: 07

Lie Groups, Chair: Adam Koranyi

H 112 Dietrich Burde Affine and projective structures on Lie groups (p.103)

Michael Wüstner

Lie groups with surjective exponential function (p.111)

Walter J. Michaelis

The exponential matrices of noncommuting matrices may commute

(p.108)

Sect.: 08

Analytic Functions I, Chair: Žarko Pavićević

H 110 Shigeyoshi Owa Some Properties for Convulutions (p.146)

Yaroslav Ya. Prytula

Generalization of Lindelöf Theorem (p.149)

Žarko Pavićević

Meromorphic Functions Generating Normal Families in an Arbitrary Open Subset of the Unit Disk, with respect to any Subgroup of the Group of Conformal Automorphisms of the Unit Disk (p.147)

Sect.: 08

Operator Algebras, Chair: Huzihiro Araki

MA 041 Grigori Amosov Cocycle conjugacy classes of Powers’ semiflow of shifts (p.113)

Alexandr V. Kosyak

On von Neumann algebras, generated by regular representations of infinitedimensional groups (p.137)

Huzihiro Araki

Factor types of infrared representations of the XY-model (p.114)

Sect.: 08

Convexity, Semigroups, and Blocking, Chair: Karl-Goswin Grosse-Erdmann

MA 309 Farhad V. Hüsseinov Closedness of convex hull and its applications (p.131)

Nobuhisa Sakakibara

Perfectness of conelike ∗-semigroups in Q^{k} (p.153)

Karl-Goswin Grosse-Erdmann

The blocking technique and the weighted Hardy inequality (p.127)

Sect.: 08

Orthogonal Polynomials and Wavelets, Chair: Thomas C. Kriecherbauer

MA 043 Mohsen Taghavi Exact values of the autocorrelation coefficients of the Rudin-Shapiro polynomials

(p.159)

Ahmed I Zayed

Wavelets in Closed Forms (p.166)

Thomas C. Kriecherbauer

Asymptotics of Orthogonal Polynomials via Integrable Methods (p.137)

Sect.: 10

Fluids, Chair: Mitsuo Morimoto

H 111 Song Jiang Large-time behavior of solutions to the equations of a one-dimensional viscous polytropic ideal gas in unbounded domains

(p.207)

Anatoli Merzon

A complete solution of the Ursell problem (p.213)

Hiroko Morimoto

On the Existence of Steady Navier-Stokes Flow in 2D Tube-like Domain under General Outflow Condition (p.214)

Sect.: 12

Particle Approximations, Stochastic Evolutions, Chair: Alfred Liemant

H 107 Terry J. Lyons Branching particle approximations to smoothing and filtering (p.259)

Krystyna Twardowska

Applications of support theorems to invariance measure theory (p.270)

Wolfgang Wagner

Stochastic particle systems approximating the stationary Boltzmann equation

(p.271)

Sect.: 12

Limit Theorems, Inequalities, and Martingales, Chair: Sergei Nagaev

H 1029 Anatoly N. Kochubei Limit Theorems for Sums of p-Adic Random Variables (p.257)

Toshitada Shintani

Uniform integrability of L^{1}-bounded matingales (p.269)

Sergei Nagaev

Probability inequalities for sums of dependent Banach space valued random variables (p.263)

Sect.: 13

Independent Sets and Coloring, Chair: Heinz Adolf Jung

H 106 Stephan Brandt Pach’s Theorem on triangle-free graphs revisited (p.274)

Kevin A. Broughan

Computing the minimal colouring of large graphs (p.274)

Ebadollah S. Mahmoodian

Defining sets in graph colorings (p.281)

Sect.: 15

Differential Algebraic Systems, Chair: Ulrich Nowak

H 2013 Rainald Ehrig Recent advances in extrapolation methods for large scale differential algebraic problems (p.299)

Friedrich Grund

Sparse-Matrix-Techniken für Parallelrechner (p.301)

Tatyana Shtykel

On the criterion of asymptotical stability for index-1 tractable DAE (p.313)

Sect.: 17

Convex Optimization, Chair: Bernd Kummer

H 2033 Anatoly Antipin To Solving of Equilibrium Programming Problems (p.335)

Evgeny G. Belousov

On Hausdorff continuity of convex polynomial mappings (p.336)

Patrick L. Combettes

Some convergence principles in convex optimization (p.338)

Poster Session: Thursday, 20 August, 14:00–15:45

Sect.: 02  Algebra

MA 144 Shavkat A. Ayupov Nilpotent Leibniz algebras of low dimension (p.13)

Valerii Vaclavovich Benyash-Krivetz

On decomposing some finitely generated groups into amalgamated free products

(p.13)

Matyas Domokos

Determinantal ideals of certain subspace arrangements (p.16)

Kiyomitsu Horiuchi An expansion of lattice (p.19)

Andrei Marcus

Derived equivalences and Dade’s conjectures (p.22)

Shigeki Matsumoto

Solutions of Sylvester Equations AX − XB = C whose coefficients matrices A and B may have common eigenvalues

(p.23)

Marek Niezgoda An unified approach to some majorization inequalities (p.25)

Yasushi Ninomiya

Group algebras and nilpotency indices of their radicals (p.25)

Carlos Renteria-Marquez

The a-invariant of some Reed-Muller codes (p.26)

Constatin I. Sonin Krull Dimension of Malcev-Neumann Rings (p.29)

Christophe Soulé

Filtrations on Higher Algebraic K-theory (p.29)

Masamichi Tokizawa

Set product residue semigroups (p.30)

Roman Zapatrin Duality and set representations for partially ordered sets (p.32)

Konstantin A. Zubrilin

The structure theory of algebras that satisfy Capelli identities (p.33)

Poster Session: Thursday, 20 August, 16:15–18:00

Sect.: 03

Number Theory and Arithmetic Algebraic Geometry

MA 144 Francis W. Clarke Closed formulas for representing primes by binary quadratic forms (p.35)

Herbert Gangl

Polylogarithmic identities in cubical higher Chow groups (p.36)

Noriaki Kimura

A simple proof of Carlitz’s congruence of Vandiver type (p.39)

Tatsuya Okada Explicit formulas for a generalized digital sum problem (p.44)

Margaret M. Robinson

Igusa local zeta functions (p.45)

Special Activities

\underline{14:00 – 18:00 H 104}

Mathematical Software Chair: Winfried Neun, ZIB, Berlin, Germany In this session, a broad spectrum of mathematical software including symbolic and numerical systems as well as interfaces and problem solutions will be presented. Details of the presentations can be found on the ICM’98 web server at URL http://elib.zib.de/ICM98/MathSoftware .

14:00 – 14:30 Ewgenij Gawrilow and Michael Joswig, Technische Universität Berlin, Germany Polymake

14:40 – 15:10 Peter Dräxler, Universität Bielefeld, Germany CREP

15:15 – 15:45 Daniel Grayson and Michael Stillman, University of Illinois, USA Macaulay 2

15:45 – 16:15 Break

16:15 – 16:45 Stephen Linton et. al., St Andrews, Scotland and Aachen, Germany GAP

16:50 – 17:20 Walter Oevel, Benno Fuchssteiner et. al., Universität Paderborn, Germany MuPAD

17:30 – 18:00 Fabrice Rouillier et. al., INRIA Lorraine, France FRISCO

Other Events

12:15  Springer Booth, Exhibition Area

Presentation of Olli Lehto’s New Book “Mathematics Without Borders, A History of the International Mathematical Union”

Reception on the occasion of the publication of this book, organized by Springer-Verlag and IMU

\underline{18:30 Urania}

Opening of the Exhibitions Mathematik zum Anfassen — Hands–on Mathematics, Mathematical Stone Sculptures, Mathematics and Ceramics Works by High–School Students

\underline{19:00 Urania}

Opening of the VideoMath Festival

\underline{19:30 Urania}

Frank Natterer, Münster, Germany The Mathematics of Imaging in Medicine and Technology